at maximum speed, an airplane traves 2700 miles against the wind in 6 hours. Flying with the wind, the plane can travel the same distance in 5 hours.
Let x be the maximum speed of the plane and y be the speed of the wind. What is the speed of the plane with no wind?

Answer by checkley71(8403)

2700/6=450 MPH IS THE SPEED AGAINST THE WIND
2700/5=540 MPH FOR THE SPEED WITH THE WIND.
(540-450)/2=90/2=45 MPH IS THE SPEED OF THE WIND.
450+45=495
540-45=495